# Artificial Intelligence NN Neural Network

### v1

@@ 1 Introduction
The responsibility of integrating into the scientific research and higher education community entails adhering to numerous behaviors and principles essential to safeguarding the integrity of educational and scientific progress. Consequently, the utilization of tools such as text editors or language enhancement applications must align with sound practices and uphold the ethical standards of scientific research and education Lund et al. (2023); Foltynek et al. (2023). Despite the substantial advancements in AI models, particularly within Natural Language Processing (NLP) applications, there remains ongoing debate about the appropriate use of these tools for text generation Leidner and Plachouras (2017); Šuster et al. (2017). This issue could significantly impact the integrity of the academic domain Tauginienė et al. (2018). Large Language Model (LLM) such as Generative Pre-trained Transformer (GPTBrown et al. (2020); OpenAI (2022, 2023), BERT Devlin et al. (2018), and XLNet Yang et al. (2020) have gained widespread acceptance among users, demonstrating such high efficiency that distinguishing between human-generated and machine-generated content has become increasingly challenging Shahid et al. (2022); Ippolito et al. (2020).
@@
In response to this surge in the use of AI techniques for text generation, various detection methods have been developed to ascertain the origin of text Solaiman et al. (2019); Fagni et al. (2021); Mitrović et al. (2023); Gehrmann et al. (2019); Mitchell et al. (2023); Su et al. (2023). These methods are categorized as supervised and unsupervised. Unsupervised methods are notable for their versatility in text detection, as they are capable of handling diverse text domains due to the nature of their pretraining  Gehrmann et al. (2019); Mitchell et al. (2023). These tools calculate probabilities and distribute them throughout the text. Under the zero-shot framework, it is presumed that AI-generated text exhibits a higher degree of probability variation compared to human-authored text, which can demonstrate both stability and variation. To assess such changes, additional text generation is required, significantly increasing execution time. The latest approach, FastDetect Bao et al. (2024), hypothesizes that AI-generated words are produced based on decisions made by the generative model, adhering to specific generation probabilities distinct from human word choices. For humans, word selection involves multiple influencing factors, rendering the output more personal and less statistically general.
@@
The majority of generated texts are fundamentally based on embedding principles in the initial stages of any Large Language Models (LLMs), where a dense vector in space is extracted for each word to encapsulate the contextual information present in the training dataset Goldberg and Levy (2014); Pennington et al. (2014). The embedding maps words or tokens into a high-dimensional continuous space, which the model employs to facilitate various stages of text generation. This embedding, coupled with additional layers and algorithms used in LLMs Vaswani et al. (2023), informs the model’s predictions for subsequent words. Understanding this generative pattern enables the identification of AI-suggested content versus natural, human-like predictions, which are not strictly bound by patterns.
@@
Among the various embedding schemes considered in this work, we place particular emphasis on the TM-based approach. The TM is an emerging machine learning technique that has demonstrated notable success in various applications Granmo (2018); Yadav et al. (2021); Berge et al. (2019); Abeyrathna et al. (2021); Maheshwari et al. (2023), including NLP and computer vision. One architecture employed within the TM framework is the Tsetlin Machine Auto-Encoder (TM-AE), which is used in NLP tasks to generate word embeddings Bhattarai et al. (2024). The TM is distinguished by its interpretability and the transparency of its output Granmo (2018); Sharma et al. (2023); Abeyrathna et al. (2023); Yadav et al. (2021). TM-AE can produce word embeddings that encapsulate the contextual information conveyed by words derived from the training dataset Kadhim et al. (2024). Consequently, TM can be leveraged to modify texts under examination, using its contextual insights to influence AI-generated text evaluation tools, thereby creating an interpretable adversarial attack that compromises their detection capabilities.
@@
2 Related Work
In related studies focusing on constructing adversarial attacks on AI-generated text detection systems, two primary approaches are identified in Huang et al. (2024) and in Sadasivan et al. (2024) respectively. The approach Huang et al. (2024) involves replacing specific tokens in the text with words generated randomly by LLMs. Multiple variants of GPT models are employed to generate a probability distribution for replacement, thereby influencing the source of the scoring. However, that method significantly increases the complexity of the attack due to the inherent opacity of LLMs, which operate as black-box models, making it challenging to rationalize token selection. Furthermore, the reliance on large-scale LLMs results in extensive execution times. Additionally, Huang et al. (2024) employs the same models utilized for scoring or generating perturbations, thereby restricting the exposure of detection models to alternative types of probability distributions. This limitation arises because these detection models are inherently familiar with the probability distributions of the tokens they analyze. Conversely, approach  Sadasivan et al. (2024) proposes relying on a single LLM as the source for token replacement, thereby creating a paraphrase-based attack. This approach is more constrained than the former, as it significantly limits the diversity of probability distributions considered.
@@ 3 AI-Text Detection Exploration
To understand the process of creating an adversarial attack on AI-text detection tools, it is essential first to comprehend the underlying mechanisms for detection. This section will explore advanced models like DetectGPT Mitchell et al. (2023) and Fast-DetectGPT Bao et al. (2024).
@@
From the above summary of the AI-text detectors, it can be inferred that determining the origin of a text fundamentally relies on assessing the discrepancy between model-suggested probabilities and the natural probabilities characteristic of human decision-making. This principle is utilized in industrial tools, such as Turnitin, to evaluate text originality. According to their website Turnitin (2024), Turnitin employs an AI model grounded in the text-generation methodology of LLMs. The text is segmented into groups of five to ten sentences, with overlapping segments to analyze contextual inclusivity. Each segment is then processed by an AI checking model, which assigns a score between 0 and 1, indicating whether the text was human-written (0) or AI-generated (1). The overall average of these scores represents the final percentage of AI-generated content in the text.
The Turnitin model relies on the GPT-3 framework for scoring. GPT-3 is trained on extensive Internet content, allowing it to generate text by predicting the most likely next words based on its training data. This prediction process is primarily governed by the transformer architecture Vaswani et al. (2023), specifically its decoder, which determines the subsequent symbol during text generation. The key components of this mechanism are:
@@ 4.2 Implementation of Tsetlin Machine
TM Granmo (2018) effectively addresses complex pattern recognition tasks by leveraging propositional conjunctives, with each literal being managed by an individual automaton. Literals represent the entire vocabulary of the input, including their negations. For instance, in a sparse dataset where rows correspond to documents and columns correspond to literals (twice the size of the vocabulary), encoding assigns a value of 1 to a column if the corresponding $word$ exists in a document, while its negation receives the opposite value. If the $word$ does not appear in the document, the value is 0, and its negation is set to 1.
@@
The TM-AE architecture Bhattarai et al. (2024) employed in this design is based on an enhanced TM structure Glimsdal and Granmo (2021), incorporating a coalesced weight matrix $W$ to facilitate the simultaneous training of multiple outputs and enabling nested elections among clauses. For example, for the word heart, the weight assigned to the clause containing love and woman may differ from the weight assigned to the clause containing old and hospital, depending on the training context. Particularly, in this work, the weights from TM-AE are utilized to compute similarity probabilities for target words.
5 Experimental Settings
A variety of datasets, Pre-Trained Language Models (PLMs), AI-text detection models, and embedding models were utilized in this study. Except for the embedding models, the remaining datasets and models align with those used in Bao et al. (2024), aiming to evaluate the proposed approach comprehensively. This diversity ensures a broad evaluation scope, encompassing various scenarios across downstream applications. The following subsections provide a concise overview of the datasets, PLMs, detection models, and embedding models employed.
5.1 Datasets
Three English datasets from diverse domains were used in the experiments. The XSum dataset, introduced by Narayan et al. (2018), is designed for abstractive summarization and provides 500 concise summaries of news articles. The SQuAD dataset Rajpurkar et al. (2016), based on Wikipedia contexts, includes 300 samples for training models to answer questions. The Writing Prompts dataset Fan et al. (2018) contains 500 samples aimed at generating creative and coherent stories based on prompts. These datasets address distinct yet overlapping text generation tasks, enhancing various natural language processing capabilities.
5.2 Pre-trained Language Models (PLMs)
5.3 AI-Text Detection Models
The study utilized a set of classifiers designed for zero-shot evaluation of adversarial attacks, including: Fast-Detect and Detect-GPT: As detailed in Section 3. Normalized Perturbation Rank (NPR) and Log Probability and Log Rank (LRR): Both leverage rank- and probability-based features to enhance accuracy while balancing computational efficiency Su et al. (2023). LogRank, Likelihood, and Rank: Metrics based on token probabilities and their ranks to evaluate the likelihood of text being AI-generated Gehrmann et al. (2019); Solaiman et al. (2019). Entropy and Divergent N-Gram Analysis (DNA): Techniques focusing on distributional irregularities and n-gram variations to detect machine-generated text Ippolito et al. (2020); Yang et al. (2023).
5.4 Embedding Models
A diverse set of embedding models was employed to calculate token probability distributions, ensuring variety in the representation of text. The embedding models used include: GloVe: Pre-trained embeddings capturing word relationships based on co-occurrence statistics. FastText: Embeddings incorporating character-level information, effectively handling out-of-vocabulary words. Word2Vec: Static embeddings generated using CBOW or Skip-Gram methods for context prediction. TM-AE: As detailed in Section 4.2, this model uses logical expressions for word embeddings. ELMo: Contextual embeddings derived from bidirectional language models (BiLSTMs). BERT: Contextual embeddings utilizing Transformers, capturing bidirectional context. These models were trained on the One Billion Word dataset Chelba et al. (2013), with a vocabulary size limited to 40,000 tokens.
@@ 6 Results
In the experiments, the Area Under the Receiver Operating Characteristic (AUROC) score was utilized to evaluate the performance of AI-generated text detection models. An AUROC score of 1.0 signifies perfect detection, indicating the model’s certainty that the text was AI-generated. Conversely, an AUROC score of 0.5 represents a random detection performance. Perturbations were generated for each dataset sample using PLM source models, as outlined in Bao et al. (2024), which served as the baseline source for text samples in this study. Subsequently, an additional perturbation was applied based on the proposed approach, as detailed below.
@@ 6.1 Experiments with Embedding Vector of Similarity
Figure 3: Heatmap illustrating the AUROC scores across various AI-text detection methods and embedding models. The x-axis represents the detection methods, while the y-axis corresponds to embedding models. Darker colors indicate higher detection performance, with a notable decline in certain methods like “Rank”, highlighting their vulnerability to adversarial perturbations.
@@
Key Findings: Fast-DetectGPT and Entropy were the most resilient detection methods, maintaining AUROC scores above 0.8 for several embedding models. Rank performed poorly, with scores dropping below 0.1 for many embedding models. GloVe achieved favorable results, effectively lowering detection performance across various methods. BERT was least effective in representing adversarial attacks, consistently achieving high detection scores, corroborating earlier findings about its alignment with expected probability distributions. TM-AE and Word2Vec delivered consistently strong performance across different detection methods, demonstrating their utility for adversarial attacks. For detailed experimental results, refer to the tables provided in the appendix, which outline the performance of each method individually.
@@ 6.2 Experiments Utilizing Synonym Similarity Vectors
Figure 4: Impact of disturbance percentage and perturbation threshold on detection accuracy. The left panel shows the effect of increasing the percentage of replaced words, while the right panel illustrates the influence of synonym proximity (min, mid, high) within the probability vector for various PLMs.
@@ 6.3 Hybrid Experiments Using Synonym and Embedding Similarity Vectors
In TM-AE, the knowledge associated with a word is derived from the documents within the training database. For example, the target word car might be represented by a clause containing engine and not wing after training. This representation can be based on a set of training documents where car frequently appears alongside engine but not with wing. Notably, during the preparation of training data, word frequency is disregarded. The model considers only the presence or absence of a word (a Boolean value) when updating the corresponding column in the vocabulary of the input sparse vector. Essentially, the TM operates like an electoral system, where clauses vote on the target word using weights that are iteratively updated during training, ultimately forming a detailed and interpretable description of the word. Further details on the interpretability of the TM model can be found in works such as Bhattarai et al. (2024); Yadav et al. (2021, 2022).
@@
The Fast-DetectGPT results (right) show average detection scores fluctuating between 0.2 and 0.5. Specifically, detection scores for XSum varied from 0.2744 to 0.3984, while scores for SQuAD ranged from 0.3532 to 0.5030. Compared to similar work Huang et al. (2024), these results demonstrate a substantial reduction in detection scores. For the XSum dataset, the detection score decreased from 0.4431 to 0.2744 (Percentage Reduction = 38.07%), highlighting a notable decline in the accuracy of text origin detection. Similarly, for the SQuAD dataset, the score dropped from 0.5068 to 0.3532 (Percentage Reduction = 30.30%), indicating the effectiveness of the proposed hybrid substitution method in successfully representing an adversarial attack.
@@ 7 Conclusion
This work inherently involves ethical considerations, as it explores methods that could potentially bypass systems designed to detect the origin of text. Such systems often play a critical role in evaluating academic, professional, or creative works, and misuse of these methods could lead to ethical challenges by undermining trust and accountability. The primary aim of this research is to advance cybersecurity by examining the vulnerabilities of text detection systems and identifying potential adversarial strategies that could compromise their reliability. Specifically, this study highlights the role of embedding models in crafting adversarial attacks and emphasizes the need for detection systems to incorporate robust measures against such vulnerabilities. To address this, we propose the development of hybrid detection systems that integrate embedding model-based probability distributions with those from large language models (LLMs), as suggested in Huang et al. (2024), to enhance their resilience against adversarial attacks.
@@
8 Detailed Experimental Results

### v2

@@ 1 Introduction
The responsibility of integrating into the scientific research and higher education community entails adhering to numerous behaviors and principles essential to safeguarding the integrity of educational and scientific progress. Consequently, the utilization of tools such as text editors or language enhancement applications must align with sound practices and uphold the ethical standards of scientific research and education (Lund et al., 2023; Foltynek et al., 2023). Despite the substantial advancements in AI models, particularly within Natural Language Processing (NLP) applications, there remains ongoing debate about the appropriate use of these tools for text generation (Leidner & Plachouras, 2017; Šuster et al., 2017). This issue could significantly impact the integrity of the academic domain (Tauginienė et al., 2018). Large Language Model (LLM) such as Generative Pre-trained Transformer (GPT(Brown et al., 2020; OpenAI, 2022; 2023), BERT (Devlin et al., 2018), and XLNet (Yang et al., 2020) have gained widespread acceptance among users, demonstrating such high efficiency that distinguishing between human-generated and machine-generated content has become increasingly challenging (Shahid et al., 2022; Ippolito et al., 2020).
@@
In response to this surge in the use of AI techniques for text generation, various detection methods have been developed to ascertain the origin of text (Solaiman et al., 2019; Fagni et al., 2021; Mitrović et al., 2023; Gehrmann et al., 2019; Mitchell et al., 2023; Su et al., 2023). These methods are categorized as supervised and unsupervised. Unsupervised methods are notable for their versatility in text detection, as they are capable of handling diverse text domains due to the nature of their pretraining  (Gehrmann et al., 2019; Mitchell et al., 2023). These tools calculate probabilities and distribute them throughout the text. Under the zero-shot framework, it is presumed that AI-generated text exhibits a higher degree of probability variation compared to human-authored text, which can demonstrate both stability and variation. To assess such changes, additional text generation is required, significantly increasing execution time. The latest approach, FastDetect (Bao et al., 2024), hypothesizes that AI-generated words are produced based on decisions made by the generative model, adhering to specific generation probabilities distinct from human word choices. For humans, word selection involves multiple influencing factors, rendering the output more personal and less statistically general.
@@
The majority of generated texts are fundamentally based on embedding principles in the initial stages of any Large Language Models (LLMs), where a dense vector in space is extracted for each word to encapsulate the contextual information present in the training dataset (Goldberg & Levy, 2014; Pennington et al., 2014). The embedding maps words or tokens into a high-dimensional continuous space, which the model employs to facilitate various stages of text generation. This embedding, coupled with additional layers and algorithms used in LLMs (Vaswani et al., 2017), informs the model’s predictions for subsequent words. Understanding this generative pattern enables the identification of AI-suggested content versus natural, human-like predictions, which are not strictly bound by patterns.
@@
Among the various embedding schemes considered in this work, we place particular emphasis on the TM-based approach. The TM is an emerging machine learning technique that has demonstrated notable success in various applications (Granmo, 2018; Yadav et al., 2021; Berge et al., 2019; Abeyrathna et al., 2021; Maheshwari et al., 2023), including NLP and computer vision. One architecture employed within the TM framework is the Tsetlin Machine Auto-Encoder (TM-AE), which is used in NLP tasks to generate word embeddings (Bhattarai et al., 2024). The TM is distinguished by its interpretability and the transparency of its output (Granmo, 2018; Sharma et al., 2023; Abeyrathna et al., 2023; Yadav et al., 2021). TM-AE can produce word embeddings that encapsulate the contextual information conveyed by words derived from the training dataset (Kadhim et al., 2024). Consequently, TM can be leveraged to modify texts under examination, using its contextual insights to influence AI-generated text evaluation tools, thereby creating an interpretable adversarial attack that compromises their detection capabilities.
@@ 2 AI-Text Detection Exploration
To understand the process of creating an adversarial attack on AI-text detection tools, it is essential first to comprehend the underlying mechanisms for detection. This section will explore advanced models like DetectGPT (Mitchell et al., 2023) and Fast-DetectGPT (Bao et al., 2024).
@@
From the above summary of the AI-text detectors, it can be inferred that determining the origin of a text fundamentally relies on assessing the discrepancy between model-suggested probabilities and the natural probabilities characteristic of human decision-making. This principle is utilized in industrial tools, such as Turnitin, to evaluate text originality. According to their website (Turnitin, 2025), Turnitin employs an AI model grounded in the text-generation methodology of LLMs. The text is segmented into groups of five to ten sentences, with overlapping segments to analyze contextual inclusivity. Each segment is then processed by an AI checking model, which assigns a score between 0 and 1, indicating whether the text was human-written (0) or AI-generated (1). The overall average of these scores represents the final percentage of AI-generated content in the text.
The Turnitin model relies on the GPT-3 framework for scoring. GPT-3 is trained on extensive Internet content, allowing it to generate text by predicting the most likely next words based on its training data. This prediction process is primarily governed by the transformer architecture (Vaswani et al., 2017), specifically its decoder, which determines the subsequent symbol during text generation. The key components of this mechanism are:
@@ 3.2 Implementation of Tsetlin Machine
TM (Granmo, 2018) effectively addresses complex pattern recognition tasks by leveraging propositional conjunctives, with each literal being managed by an individual automaton. Literals represent the entire vocabulary of the input, including their negations. For instance, in a sparse dataset where rows correspond to documents and columns correspond to literals (twice the size of the vocabulary), encoding assigns a value of 1 to a column if the corresponding $word$ exists in a document, while its negation receives the opposite value. If the $word$ does not appear in the document, the value is 0, and its negation is set to 1.
@@
The TM-AE architecture (Bhattarai et al., 2024) employed in this design is based on an enhanced TM structure (Glimsdal & Granmo, 2021), incorporating a coalesced weight matrix $W$ to facilitate the simultaneous training of multiple outputs and enabling nested elections among clauses. For example, for the word heart, the weight assigned to the clause containing love and woman may differ from the weight assigned to the clause containing old and hospital, depending on the training context. Particularly, in this work, the weights from TM-AE are utilized to compute similarity probabilities for target words.
@@ 4 Results
Section B in the appendix includes further details that offer a concise overview of the datasets, Pre-Trained Language Models (PLMs), detection models, and embedding models used. In the experiments, the Area Under the Receiver Operating Characteristic (AUROC) score was utilized to evaluate the performance of AI-generated text detection models. An AUROC score of 1.0 signifies perfect detection, indicating the model’s certainty that the text was AI-generated. Conversely, an AUROC score of 0.5 represents a random detection performance. Perturbations were generated for each dataset sample using PLM source models, as outlined in Bao et al. (2024), which served as the baseline source for text samples in this study. Subsequently, an additional perturbation was applied based on the proposed approach, as detailed below.
@@ 4.1 Experiments with Embedding Vector of Similarity
Figure 3: Heatmap illustrating the AUROC scores across various AI-text detection methods and embedding models. The x-axis represents the detection methods, while the y-axis corresponds to embedding models.
@@
Key Findings: Fast-DetectGPT and Entropy were the most resilient detection methods, maintaining AUROC scores above 0.8 for several embedding models. Rank performed poorly, with scores dropping below 0.1 for many embedding models. GloVe achieved favorable results, effectively lowering detection performance across various methods. BERT was least effective in representing adversarial attacks, consistently achieving high detection scores, corroborating earlier findings about its alignment with expected probability distributions. TM-AE and Word2Vec delivered consistently strong performance across different detection methods, demonstrating their utility for adversarial attacks. For detailed experimental results, refer to the tables provided in the appendix (Section C), which outline the performance of each method individually.
@@ 4.2 Experiments Utilizing Synonym Similarity Vectors
Figure 4: Impact of disturbance percentage and perturbation threshold on detection accuracy. The left panel shows the effect of increasing the percentage of replaced words, while the right panel illustrates the influence of synonym proximity (min, mid, high).
@@ 4.3 Hybrid Experiments Using Synonym and Embedding Similarity Vectors
In TM-AE, the knowledge associated with a word is derived from the documents within the training database. For example, the target word car might be represented by a clause containing engine and not wing after training. This representation can be based on a set of training documents where car frequently appears alongside engine but not with wing. Notably, during the preparation of training data, word frequency is disregarded. The model considers only the presence or absence of a word (a Boolean value) when updating the corresponding column in the vocabulary of the input sparse vector. Essentially, the TM operates like an electoral system, where clauses vote on the target word using weights that are iteratively updated during training, ultimately forming a detailed and interpretable description of the word. Further details on the interpretability of the TM model can be found in works such as Bhattarai et al. (2024),  Yadav et al. (2021), and Yadav et al. (2022).
@@
The Fast-DetectGPT results (right) show average detection scores fluctuating between 0.2 and 0.5. Specifically, detection scores for XSum varied from 0.2744 to 0.3984, while scores for SQuAD ranged from 0.3532 to 0.5030. Compared to similar work (Huang et al., 2024), these results demonstrate a substantial reduction in detection scores. For the XSum dataset, the detection score decreased from 0.4431 to 0.2744 (Percentage Reduction = 38.07%), highlighting a notable decline in the accuracy of text origin detection. Similarly, for the SQuAD dataset, the score dropped from 0.5068 to 0.3532 (Percentage Reduction = 30.30%), indicating the effectiveness of the proposed hybrid substitution method in successfully representing an adversarial attack.
@@ 5 Conclusion
This work inherently involves ethical considerations, as it explores methods that could potentially bypass systems designed to detect the origin of text. Such systems often play a critical role in evaluating academic, professional, or creative works, and misuse of these methods could lead to ethical challenges by undermining trust and accountability. The primary aim of this research is to advance cybersecurity by examining the vulnerabilities of text detection systems and identifying potential adversarial strategies that could compromise their reliability. Specifically, this study highlights the role of embedding models in crafting adversarial attacks and emphasizes the need for detection systems to incorporate robust measures against such vulnerabilities. To address this, we propose the development of hybrid detection systems that integrate embedding model-based probability distributions with those from large language models, as suggested in Huang et al. (2024), to enhance their resilience against adversarial attacks.
@@
Appendix A Related Work
In related studies focusing on constructing adversarial attacks on AI-generated text detection systems, two primary approaches are identified in Huang et al. (2024) and in Sadasivan et al. (2024), respectively. The approach Huang et al. (2024) involves replacing specific tokens in the text with words generated randomly by LLMs. Multiple variants of GPT models are employed to generate a probability distribution for replacement, thereby influencing the source of the scoring. However, that method significantly increases the complexity of the attack due to the inherent opacity of LLMs, which operate as black-box models, making it challenging to rationalize token selection. Furthermore, the reliance on large-scale LLMs results in extensive execution times. Additionally, Huang et al. (2024) employs the same models utilized for scoring or generating perturbations, thereby restricting the exposure of detection models to alternative types of probability distributions. This limitation arises because these detection models are inherently familiar with the probability distributions of the tokens they analyze. Conversely, approach  Sadasivan et al. (2024) proposes relying on a single LLM as the source for token replacement, thereby creating a paraphrase-based attack. This approach is more constrained than the former, as it significantly limits the diversity of probability distributions considered.
Appendix B Experimental Resources
A variety of datasets, PLMs, AI-text detection models, and embedding models were utilized in this study. Except for the embedding models, the remaining datasets and models align with those used in Bao et al. (2024), aiming to evaluate the proposed approach comprehensively. This diversity ensures a broad evaluation scope, encompassing various scenarios across downstream applications. The following subsections provide a concise overview of the datasets, PLMs, detection models, and embedding models employed.
B.1 Datasets
Three English datasets from diverse domains were used in the experiments. The XSum dataset, introduced by Narayan et al. (2018), is designed for abstractive summarization and provides 500 concise summaries of news articles. The SQuAD dataset (Rajpurkar et al., 2016), based on Wikipedia contexts, includes 300 samples for training models to answer questions. The Writing Prompts dataset (Fan et al., 2018) contains 500 samples aimed at generating creative and coherent stories based on prompts. These datasets address distinct yet overlapping text generation tasks, enhancing various natural language processing capabilities.
B.2 Pre-trained Language Models (PLMs)
B.3 AI-Text Detection Models
The study utilized a set of classifiers designed for zero-shot evaluation of adversarial attacks, including: Fast-Detect and Detect-GPT: As detailed in Section 2. Normalized Perturbation Rank (NPR) and Log Probability and Log Rank (LRR): Both leverage rank- and probability-based features to enhance accuracy while balancing computational efficiency (Su et al., 2023). LogRank, Likelihood, and Rank: Metrics based on token probabilities and their ranks to evaluate the likelihood of text being AI-generated (Gehrmann et al., 2019; Solaiman et al., 2019). Entropy and Divergent N-Gram Analysis (DNA): Techniques focusing on distributional irregularities and n-gram variations to detect machine-generated text (Ippolito et al., 2020; Yang et al., 2023).
B.4 Embedding Models
A diverse set of embedding models was employed to calculate token probability distributions, ensuring variety in the representation of text. The embedding models used include: GloVe: Pre-trained embeddings capturing word relationships based on co-occurrence statistics. FastText: Embeddings incorporating character-level information, effectively handling out-of-vocabulary words. Word2Vec: Static embeddings generated using CBOW or Skip-Gram methods for context prediction. TM-AE: As detailed in Section 3.2, this model uses logical expressions for word embeddings. ELMo: Contextual embeddings derived from bidirectional language models (BiLSTMs). BERT: Contextual embeddings utilizing Transformers, capturing bidirectional context. These models were trained on the One Billion Word dataset (Chelba et al., 2013), with a vocabulary size limited to 40,000 tokens.
Appendix C Detailed Experimental Results